zarkam21 (zarkam21):

If the bases of an isosceles trapezoid have lengths of 15 and 32, what is the length of the median. If necessary, round your answer to two decimal places.

OpenStudy (shelbywyatt):

A B + C D 2

OpenStudy (shelbywyatt):

ab+cd/2

OpenStudy (shelbywyatt):

15+32=47/2=?

zarkam21 (zarkam21):

23.5

OpenStudy (shelbywyatt):

yeah